OpenStudy (anonymous):

Someone help me simplify an expression! Someone help me simplify an expression! @Mathematics

OpenStudy (sasogeek):

what expression...

OpenStudy (anonymous):

\[\sqrt[4]{36}\times \sqrt[4]{9}\div \sqrt[4]{4}\]

OpenStudy (anonymous):

the Answer is 3

OpenStudy (anonymous):

Thanks! :D

OpenStudy (sasogeek):

yw